# Do Language Models Have a Critical Period for Language Acquisition?

### v1

@@ 4 Research Questions and Methodology
These are the main questions that we want to investigate, and on which we focus our experiments. We analyse them by first training LMs in various multilingual setups—while altering the ages at which $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ are acquired—and then evaluating our models on $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$. Importantly, we do not make any modifications to the LMs’ architecture or learning objectives in these experiments.
@@
Investigating RQ 3 serves a dual purpose. First, it allows us to test whether the critical period effects associated with $\mathrm{L}_{2}$ acquisition and $\mathrm{L}_{1}$ attrition arise in tandem as a result of a loss in plasticity, or whether these phenomena can (at least in one case) be decoupled. Second, if we are successful at reverse engineering these CP effects, we obtain a model which more closely resemble human learners; and this might be useful for future work. We test this question by evaluating the linguistic performance on both $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ while training a model whose learning objective has an extra regularizer which enforces a reduction in plasticity.
@@ 4.1 Training conditions
Our goal is to train language models $\mathbb{L}$ from scratch on pairs of languages ($\mathrm{L}_{1}$ and $\mathrm{L}_{2}$) while manipulating two independent variables: (i) the ages of exposure to $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$, and (ii) the level of programmed plasticity. In this section our focus will be on providing the methodology for variable (i), whereas the methodology for variable (ii) is left to be presented in Section 4.2.
The obvious way to manipulate age of acquisition in the case of language modeling is to alter the training data schedule. We derive five such scheduling layouts, to which we will refer as “training conditions” throughout the paper. A visualization of these five conditions is provided in Fig. 1. Across all conditions, the datasets remain unchanged and the size of the training data per language is kept consistent (we denote this quantity by $\mathrm{S}$). The total number of training iterations per example, also known as epochs, is also a constant number (we refer to this as $\mathrm{E}$). As a consequence, the amount of exposure to $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ is the same across conditions, with the inevitable exceptions of the monolingual and sequential-interleaved conditions.
@@
This condition seeks to imitate a monolingual human learner who is exposed to only one language during his lifetime. The simplest approach in this condition would be to just train $\mathbb{L}$ for $2\cdot\mathrm{E}$ epochs on a monolingual dataset. However, this would mean that $\mathbb{L}$ would be trained only on half the amount of tokens ($\mathrm{S}$) compared to the other conditions ($2\cdot\mathrm{S}$). To account for this, we create a second monolingual dataset of the same size and train $\mathbb{L}$ on the two datasets in a sequential manner.
@@
This condition aims to replicate a bilingual human learner that is exposed to two different languages at the same time. In the context of CP we consider this a simulation of a young $\mathrm{L}_{2}$ learner. We achieve this by training $\mathbb{L}$ for a total of $\mathrm{E}$ epochs on an interleaved bilingual dataset. Throughout training, $\mathbb{L}$ encounters batches of fixed size that alternate between $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$. As the bilingual dataset is double in size compared to the monolingual datasets, $\mathrm{E}$ epochs provide the same amount of training steps per language as with the other conditions.
@@
This condition represents a straightforward way to reproduce the idea of sequential learning ($\mathrm{L}_{1}$ followed by $\mathrm{L}_{2}$). In the context of CP we consider this a simulation of an adult $\mathrm{L}_{2}$ learner. In this condition, $\mathbb{L}$ is trained for $\mathrm{E}$ epochs exclusively on $\mathrm{L}_{1}$, and then subsequently trained for another $\mathrm{E}$ epochs on $\mathrm{L}_{2}$. It is worth noting that the shift from $\mathrm{L}_{1}$ to $\mathrm{L}_{2}$ occurs abruptly, with a complete halt on $\mathrm{L}_{1}$ exposure rather than a gradual transition.
@@
This condition is closely related to the previous sequential condition, but recreates a human-like continual learning environment. Typically, when acquiring their second language, individuals continue to engage with their first language (although there may be exceptions to this pattern). This condition continues $\mathrm{L}_{1}$ exposure interleaved with $\mathrm{L}_{2}$ in the second stage of the training process. We continue to use the same $\mathrm{L}_{1}$ dataset from the initial training stage.
@@
This condition attempts to emulate a reduction in plasticity in LMs. The $\mathrm{L}_{2}$ models are trained from the same $\mathrm{L}_{1}$ checkpoints as for the normal sequential condition, but with EWC regularization added to the loss function. The reduction in plasticity is not progressive, but only changes once, after $\mathrm{L}_{1}$ has been fully trained.
@@ 4.2 Enforcing Plasticity
There are several methods to simulate a computational reduction in plasticity in LMs. We have chosen Elastic Weight Consolidation (EWC; Kirkpatrick et al., 2017). In short, EWC introduces a Bayesian-inspired regularization term on a LM’s loss partway through training; this term penalizes deviations from a prior distribution over the parameter space (defined in terms of $\mathrm{L}_{1}$ training), simulating the end of the critical period. The modified loss function is defined as
@@
where the second term corresponds to the well-known L2 regularization that results from the prior over the model parameters $\boldsymbol{\theta}$. We introduce an additional parameter $\lambda\in\mathbb{R}_{\geq 0}$ to control the strength of the EWC regularization term. In the trivial case when $\lambda=0$, there is no programmed decrease in plasticity. We provide the complete derivation of EWC in App. E, including the necessary approximations we make for its practical implementation.
5 Experimental Setup
This section provides a comprehensive description of the experimental setup for this project, including: languages, dataset, model architecture, and evaluation methods.
@@ 5 Experimental Setup
Nonetheless, to improve the generalizability of our results, we also run one experiment with an extended set of languages (see Experiment 4 in § 6). On that account, a more diversified set of languages varying language family (Indo-European (IE) or not) and script (Latin (L) or not) is selected as follows (from most to least related):
@@
We introduce a special category labeled “programming languages” as an extra baseline. We classify these languages as having notably distinct structural and syntactic properties, potentially leading to different learning behaviours.
@@
It is desirable to acquire data that closely resembles the kinds of language children encounter during their formative years. This typically involves natural speech and narratives. Unfortunately, there are no existing datasets in any language that are fully representative of the type and volume of language input a child is exposed to during learning, let alone in a variety of languages. As a consequence, in this work, we construct a customized mix of multilingual training data sourced from three complementary domains: spoken, literature and non-fiction. We select the OpenSubtitles (Lison and Tiedemann, 2016) corpus for the spoken domain, the Gutenberg (Gerlach and Font-Clos, 2018) collection for literature, and the Wikipedia content for non-fiction.666 To perform Experiment 4 with a broader range of languages, we exclude the Gutenberg dataset due to insufficient data availability across languages. Further, for the special Java (programming) language, we use an additional corpus called The Stack (Kocetkov et al., 2022). While these choices of datasets are not fully developmentally plausible, we judge them to have a good balance of diversity, quality, and quantity among the limited set of publicly available multilingual datasets. Data preprocessing details can be found in App. A and statistics about the data are in App. B.
@@
We run our experiments with both autoregressive and masked language models. To represent these two categories, we consider RoBERTa (Liu et al., 2019), a encoder-only transformer trained to predict masked tokens, and GPT-2 (Radford et al., 2019), a decoder-only transformer trained to predict next tokens. We rely on implementations from the HuggingFace Transformers library (Wolf et al., 2020), namely roberta-base with 125M parameters and gpt2 with 137M parameters. We choose these two models due to their availability, size and scientific relevance. These models are then trained according to the conditions defined in § 4.1. In all three sequential conditions, $\mathrm{L}_{2}$ training starts from the final $\mathrm{L}_{1}$ checkpoints, but with a new optimizer.
@@
We run a Bayesian hyperparameter search using the Weights & Biases Sweeps API (Biewald, 2020) to identify good model configurations for our training data and methodology. From the sweep results, we extract several model configurations (given in App. C) with which we run our various experiments in § 6.
@@
We use PPL, BLiMP and GLUE to assess models’ language proficiency throughout our experiments. We note that both BLiMP and GLUE  [^8] We use GLUE to refer to a combination of tasks from GLUE and superGLUE (see Appendix D) implemented in the BabyLM evaluation pipeline (Warstadt et al., 2023). are exclusively available in English. Throughout this work, we thus use English as the language in which we perform evaluations. While this is far from optimal from the perspective of linguistic representation, we prioritize fine-grained evaluation here, rather than testing a breadth of languages using only PPL.
@@ 6 Experiments
This section describes our four experiments, designed to provide sufficient supporting evidence for the research questions detailed in § 4.
@@
Figure 2: $\mathrm{L}_{2}$ ($\mathtt{en}$) results for regular training (6 epochs). Results are aggregated across model configuration and $\mathrm{L}_{1}$ ($\mathtt{de}$ and $\mathtt{fi}$). Top: PPL on $\mathrm{L}_{2}$ ($\mathtt{en}$) during training on $\mathrm{L}_{2}$. Middle: Accuracy on BLiMP during training on $\mathrm{L}_{2}$. Bottom: Accuracy on GLUE at the end of training.
@@
The goal of this setup is to study the CP for $\mathrm{L}_{2}$ learning, which relates to RQ 1 and RQ 3. We run this experiment on both GPT-2 and RoBERTa models, using German and Finnish data as a first language and English data as a second language. The dataset for each language has the same size of 600 million tokens, a factor which was limited by the availability of the Finnish data. The models are trained with a limited budget of 6 epochs per language. This number has been empirically determined (after preliminary exploration) to provide a good trade-off between computational costs and model performance (i.e. the models perform sufficiently well after training for 6 epochs and the learning slows down). Finally, in order to introduce more variability and provide more result samples, the trainings are run with three different configurations ($\mathbf{C_{1},C_{2},C_{3}}$, see App. C).
We illustrate the results for this experiment in Fig. 2. In general, the learning patterns through epochs are similar across conditions (except for interleaved), with variations observed mostly in the final performance. As expected, the monolingual performance is the best among all conditions. [^9] The monolingual PPL is slightly increased artificially due to the usage of a monolingual tokenizer instead of a bilingual. Given a vocabulary size, the bilingual tokenizer will have fewer English-specific tokens and more tokens per word on average. Hence we should rely on the other two metrics for comparisons between bilingual and monolingual conditions. We notice that the interleaved condition performs slightly worse than the sequential condition, although the difference is more noticeable in RoBERTa models than in GPT-2 models. Both achieve lower scores than the monolingual training (the baseline for native level language proficiency). Apart from this, the sequential-interleaved condition presents different outcomes based on the model architecture. On GPT-2 it is worse (i.e., keeping $\mathrm{L}_{1}$ exposure hinders $\mathrm{L}_{2}$ learning), while on RoBERTa it is better (i.e., keeping $\mathrm{L}_{1}$ exposure helps $\mathrm{L}_{2}$ learning). Lastly, the sequential-ewc condition shows clear differences in the learning behaviour, with much worse $\mathrm{L}_{2}$ proficiency. This means that the reduction in plasticity through regularization has a significant (negative) impact on $\mathrm{L}_{2}$’s learning outcome.
@@
Figure 3: $\mathrm{L}_{1}$ ($\mathtt{en}$) results when the language order is reversed (6 + 6 epochs). Results are aggregated across $\mathrm{L}_{2}$ ($\mathtt{de}$ and $\mathtt{fi}$). Top: PPL (log scale) on the $\mathrm{L}_{1}$ ($\mathtt{en}$) validation set during training. Middle: Accuracy on BLiMP during training. Bottom: Accuracy on GLUE at the end of training on $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$.
@@
The results for this experiment are displayed in Fig. 3. In general, it can be noticed that there is a smaller drop (from epoch 6 to 12) in GLUE scores compared to BLiMP scores. This is most probably caused by the conceptual difference between evaluations in a zero-shot setting and evaluations that require fine-tuning. When fine-tuning, LMs are still able to learn to solve the new downstream tasks, even though the foundational language knowledge is worse. In the monolingual condition, $\mathrm{L}_{1}$ performance keeps improving until the end of training, even though it slows down in the second stage. In the interleaved condition, models reach the same level of $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ proficiency (when comparing these results to the ones from Fig. 2). The more notable result comes from the sequential learning. In this condition, LMs start to lose the knowledge acquired from $\mathrm{L}_{1}$ learning right after $\mathrm{L}_{2}$ exposure is started. The final $\mathrm{L}_{1}$ perplexity values indicate that, in the end, the LMs forget almost everything that they have learned before. It looks like the second language completely replaces the first one. Finally, an interesting shift in behavior occurs within the sequential-interleaved condition. When $\mathrm{L}_{1}$ exposure is prolonged without any reduction in plasticity, models are able to retain all the prior $\mathrm{L}_{1}$ knowledge. [^10] RoBERTa continues learning $\mathrm{L}_{1}$ during the second stage. We believe this is due to the fact that RoBERTa models are undertrained in comparison with GPT-2.
Figure 4: Summary of the $\mathrm{L}_{2}$ ($\mathtt{en}$) evaluation results for the convergence training (48 epochs). Results are aggregated across $\mathrm{L}_{1}$ ($\mathtt{de}$ and $\mathtt{fi}$). Top left: PPL (log scale) on the $\mathrm{L}_{2}$ ($\mathtt{en}$) validation set during training on $\mathrm{L}_{2}$. Top right: PPL (log scale) on the $\mathrm{L}_{1}$ ($\mathtt{de}$, $\mathtt{fi}$) validation set during training. Bottom left: Accuracy on BLiMP during $\mathrm{L}_{2}$ training. Bottom right: Accuracy on GLUE at the end of $\mathrm{L}_{2}$ training.
@@
However, as we have seen from Fig. 2, $\mathrm{L}_{2}$ learning is also harmed in this case. To explore this trade-off, we vary the $\lambda$ values and test the models’ performance on both $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ at the end of training. The results are illustrated in Fig. 5. We see that the regularization strength $\lambda$ highly influences both $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ learning outcomes. When high $\lambda$ values are used (strong EWC regularization), $\mathrm{L}_{1}$ knowledge can be predictably maintained at the initial levels. However, $\mathrm{L}_{2}$ learning will be almost completely impaired. On the other side, when lower $\lambda$ values are used (weak EWC regularization), $\mathrm{L}_{2}$ learning is affected less, but $\mathrm{L}_{1}$ will not be preserved. It is also noticeable that in this case, there is a higher variance on the final $\mathrm{L}_{1}$ outcomes. As mentioned in § 5, we have selected the $\lambda$ values for all our experiments as the point where $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ performance is roughly equivalent (intersection lines are marked on the plots). This is such that we do not favor either very strong or very weak regularization, and also to match the behaviour exposed by the interleaved models.
Figure 6: $\mathrm{L}_{2}$ ($\mathtt{en}$) results for each language pair at the end of training (6 epochs). Models are trained using the GPT-2 architecture. We include the score of the $\mathtt{en}$ monolingual baseline ($\mathrm{L}_{1}$ = $\mathtt{en}2$) in each condition for direct comparison. Top: Perplexity on $\mathrm{L}_{2}$ ($\mathtt{en}$) at the end of training. Bottom: Accuracy on BLiMP at the end of training.
@@
This experiment is motivated by the observation that CP effects can become stronger with a later age of exposure. We extend the training time for each language (thus postponing the start of $\mathrm{L}_{2}$ exposure) to 48 epochs, allowing the model weights to better converge. As this is more computationally demanding, we do only one run per condition.
The results are provided in Fig. 4. This figure shows that the final results are more uniform across the first three conditions, especially between interleaved and sequential. The performance of the interleaved setting shows a slight improvement compared to regular training, indicating that a longer training duration was beneficial (and necessary for convergence). Notably, the learning pattern for the sequential-ewc condition does not change, the final $\mathrm{L}_{2}$ performance does not significantly improve with longer training. It looks like the EWC regularization does not simply slow down $\mathrm{L}_{2}$ learning, but rather acts as a lower bound: $\mathrm{L}_{2}$ knowledge can never improve past a certain point for a given choice of $\lambda$.
@@
This experiment both addresses RQ 4 and provides more diversity in the results; the latter is especially important considering the limited selection of languages for the previous experiments. The main focus here is to find whether language, or more concretely, the degree of similarity between $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$, has any impact on the CP. For this, we consider a wider range of languages based on their relatedness to English (see § 5). To accommodate for the significant increase in computational demand, we consider only one model architecture in this experiment (GPT-2, config. $\mathbf{C_{5}}$), and we reduce the size of the training data to 100 million tokens.
We present the findings in Fig. 6. Our initial expectations were guided by Papadimitriou and Jurafsky’s (2020) study, who showed that the performance of LSTMs on $\mathrm{L}_{2}$ is highly correlated with the language’s syntactic similarity to $\mathrm{L}_{1}$. However, in their experiments the model weights are frozen after the initial $\mathrm{L}_{1}$ training, and only the final embedding representations are trained on $\mathrm{L}_{2}$. We do not see the same result when a full $\mathrm{L}_{2}$ fine-tuning is performed: our $\mathrm{L}_{2}$ ($\mathtt{en}$) models perform equally well regardless of the choice of $\mathrm{L}_{1}$.

### v2

@@ 4 Research Questions and Methodology
These are the main questions that we want to investigate, and on which we focus our experiments. We analyse them by first training LMs in various multilingual setups—while altering the ages (epochs) at which $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ are acquired—and then evaluating our models on $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$. Importantly, we do not make any modifications to the LMs’ architecture or learning objectives in these experiments.
@@
Investigating RQ 3 serves a dual purpose. First, it allows us to test whether the critical period effects associated with $\mathrm{L}_{2}$ acquisition and $\mathrm{L}_{1}$ attrition arise in tandem as a result of a loss in plasticity, or whether these phenomena can (at least in one case) be decoupled. Second, if we are successful at reverse engineering these CP effects, we obtain a model that more closely resembles human learners and that might be useful for future work. We test this question by evaluating the linguistic performance on both $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ while training a model whose learning objective has an extra regularizer which enforces a reduction in plasticity.
@@ 4.1 Training conditions
Our goal is to train a language model $\mathbb{L}$ from scratch on pairs of languages ($\mathrm{L}_{1}$ and $\mathrm{L}_{2}$) while manipulating two independent variables: (i) the ages (epochs) of exposure to $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$, and (ii) the level of programmed plasticity. In this section our focus will be on providing the methodology for variable (i), whereas the methodology for variable (ii) is left to be presented in Section 4.2.
The obvious way to manipulate age of acquisition in the case of language modeling is to alter the training data schedule. As visualized in Fig. 1, we consider five schedules, which we will refer to as “training conditions” throughout the paper. Across all conditions, the datasets remain unchanged (for a given language pair) and the size of the training data per language is kept consistent (we denote this quantity by $\mathrm{S}$). The total number of training iterations per example, also known as epochs, is also a constant number (we refer to this as $\mathrm{E}$). As a consequence, the amount of exposure to $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ is the same across conditions, with the inevitable exceptions of the monolingual and sequential-interleaved conditions.
@@
This condition simulates a monolingual human learner exposed to only one language during their lifetime. The simplest approach in this condition would be to just train $\mathbb{L}$ for $2\cdot\mathrm{E}$ epochs on a monolingual dataset. However, this would mean that $\mathbb{L}$ would be trained only on half the number of tokens ($\mathrm{S}$) compared to the other conditions ($2\cdot\mathrm{S}$). To account for this, we create a second monolingual dataset of the same size and train $\mathbb{L}$ on the two datasets in a sequential manner.
@@
This condition aims to replicate a simultaneous bilingual human learner exposed to two different languages from birth. It is also an implementation of typical multitask learning. We train $\mathbb{L}$ for a total of $\mathrm{E}$ epochs on an interleaved bilingual dataset. Throughout training, $\mathbb{L}$ encounters batches of fixed size that alternate between $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$. As the bilingual dataset is double in size compared to the monolingual datasets, $\mathrm{E}$ epochs provide the same amount of training steps per language as with the other conditions.
@@
This condition represents the experience of a late $\mathrm{L}_{2}$ learner who changes linguistic communities, losing exposure to $\mathrm{L}_{1}$ entirely. It can also be seen as a typical implementation of transfer learning. In this condition, $\mathbb{L}$ is trained for $\mathrm{E}$ epochs exclusively on $\mathrm{L}_{1}$, and then subsequently trained for another $\mathrm{E}$ epochs on $\mathrm{L}_{2}$. The shift from $\mathrm{L}_{1}$ to $\mathrm{L}_{2}$ occurs abruptly, with a complete halt in $\mathrm{L}_{1}$ exposure rather than a gradual transition.
@@
This condition is closely related to the previous sequential condition, but recreates an experience more common among human bilinguals where $\mathrm{L}_{1}$ exposure continues during $\mathrm{L}_{2}$ acquisition. It is also an implementation of one approach to continual learning. After the initial stage of $\mathrm{L}_{1}$ learning, $\mathbb{L}$ is trained on $\mathrm{L}_{1}$ interleaved with $\mathrm{L}_{2}$. We continue to use the same $\mathrm{L}_{1}$ dataset from the initial training stage.
@@
This condition attempts to emulate in LMs an innate reduction in plasticity like that proposed for humans. The $\mathrm{L}_{2}$ models are trained from the same $\mathrm{L}_{1}$ checkpoints as for the normal sequential condition, but with EWC regularization added to the loss function. The reduction in plasticity is not progressive, but only changes once, after $\mathrm{L}_{1}$ has been fully trained.
@@ 4.2 Enforcing Plasticity
There are several methods to simulate a computational reduction in plasticity in LMs. We have chosen Elastic Weight Consolidation (Kirkpatrick et al., 2017) due to its popularity and simplicity. EWC introduces a Bayesian-inspired regularization term on a LM’s loss partway through training; this term penalizes deviations from a prior distribution over the parameter space (defined in terms of $\mathrm{L}_{1}$ training), simulating the end of the critical period. The modified loss function is defined as
@@
We introduce an additional hyperparameter $\lambda\in\mathbb{R}_{\geq 0}$ to control the strength of the EWC regularization term. In the trivial case when $\lambda=0$, there is no programmed decrease in plasticity. We provide the complete derivation of EWC in App. E.
5 Experimental Setup [^7] The code is released at https://github.com/iconstantinescu/lm-critical-period.
This section provides a comprehensive description of the experimental setup for this project, including: languages, datasets, model architectures, and evaluation methods.
@@
Nonetheless, to improve the generalizability of our results, we also run one experiment with an extended set of languages as $\mathrm{L}_{1}$, using English as $\mathrm{L}_{2}$ (see Experiment 4 in § 6). We select languages from various language families (Indo-European (IE) or not) using different scripts (Latin (L) or not). To represent the extreme endpoints, we also use a different corpus of English as the most closely related $\mathrm{L}_{1}$, and a corpus in a programming language (Java) as the least closely related $\mathrm{L}_{1}$. The complete list is as follows (from most to least related):
  • Same language: English2 ($\mathtt{en2}$)
@@
Ideally, we would train on data that closely resembles the kinds of language children encounter during language acquisition. This typically involves natural speech and narratives. Unfortunately, there are no existing datasets in any language that are fully representative of the type and volume of language input a child is exposed to during learning, let alone in a variety of languages. As a consequence, in this work, we construct a customized mix of multilingual training data sourced from three complementary domains: spoken, literature and non-fiction. We select the OpenSubtitles (Lison and Tiedemann, 2016) corpus for the spoken domain, the Gutenberg (Gerlach and Font-Clos, 2020) collection for literature, and the Wikipedia content for non-fiction.999 To perform Experiment 4 with a broader range of languages, we exclude the Gutenberg dataset due to insufficient data availability across languages. Further, for the special Java (programming) language, we use an additional corpus called The Stack (Kocetkov et al., 2023). While these choices of datasets are not fully developmentally plausible, we judge them to have a good balance of diversity, quality, and quantity among the limited set of publicly available multilingual datasets. See App. A for preprocessing details and App. B for data statistics.
@@
We run our experiments with both autoregressive and masked language models. To represent these two categories, we consider RoBERTa (Liu et al., 2019), an encoder-only transformer trained to predict masked tokens, and GPT-2 (Radford et al., 2019), a decoder-only transformer trained to predict next tokens. We rely on implementations from the HuggingFace Transformers library (Wolf et al., 2020), namely roberta-base with 125M parameters and gpt2 with 137M parameters. We choose these two models due to their availability, size and scientific relevance. These models are then trained according to the conditions defined in § 4.1. In all three sequential conditions, $\mathrm{L}_{2}$ training starts from the final $\mathrm{L}_{1}$ checkpoints, but with a new optimizer.
@@
We run a Bayesian hyperparameter search using the Weights & Biases Sweeps API (Biewald, 2020) to identify good model configurations for our training data and methodology. We extract several model configurations (see App. C) with which we run the various experiments from § 6.
@@
We use Perplexity (PPL), BLiMP and GLUE to assess models’ language proficiency throughout our experiments. To account for the differences in tokenization across language pairs, we report the PPL per character. [^11] In addition, PPL per character being smaller in magnitude than the more familiar PPL per token, differences between PPL scores relative to the total magnitude are smaller as well. BLiMP (Warstadt et al., 2020) is a dataset of minimal pairs targeting contrasts in acceptability for a variety of grammatical phenomena in English. Evaluation is performed in a zero-shot setting by comparing LM surprisal on a grammatical and ungrammatical sentence. GLUE and superGLUE [^12] We henceforth use GLUE to refer to a combination of tasks from GLUE and superGLUE (see App. D) implemented in the BabyLM evaluation pipeline (Warstadt et al., 2023). are compilations of semantic, commonsense, and syntactic tasks (Wang et al., 2018, 2019a). Evaluation is performed by fine-tuning all parameters of the model as well as a randomly initialized classifier MLP. While similar benchmarks (some even using the names BLiMP and GLUE) are available for many of the languages we study, these are inherently different datasets which are not mutually comparable. Thus, throughout this work, we use only English as the evaluation language. This is far from optimal from the perspective of linguistic representation, but we prioritize controlled evaluation, rather than obtaining a large set of exploratory results.
@@ 6 Experiments
This section describes our four experiments, designed to provide evidence for answering the research questions detailed in § 4.
@@
Figure 2: $\mathrm{L}_{2}$ ($\mathtt{en}$) results for regular training (6 epochs). Results are aggregated across model configuration and $\mathrm{L}_{1}$ ($\mathtt{de}$ and $\mathtt{fi}$). Top: PPL per character on $\mathrm{L}_{2}$ ($\mathtt{en}$) during training on $\mathrm{L}_{2}$. Middle: Accuracy on BLiMP during training on $\mathrm{L}_{2}$. Bottom: Performance on GLUE at the end of training.
@@
The goal of this setup is to study the CP for $\mathrm{L}_{2}$ learning, which relates to RQ 1 and RQ 3. We run this experiment on both GPT-2 and RoBERTa models, using German and Finnish data as a first language and English data as a second language. The dataset for each language has the same size of 600 million tokens, a factor which was limited by the availability of the Finnish data. The models are trained with a limited budget of 6 epochs per language. This number has been empirically determined (after preliminary exploration) to provide a good trade-off between computational costs and model performance (i.e., the models perform sufficiently well after training for 6 epochs and the learning slows down). Finally, in order to introduce more variability and provide more result samples, the trainings are run with three different configurations ($\mathbf{C_{1},C_{2},C_{3}}$, see App. C).
@@
Figure 3: $\mathrm{L}_{1}$ ($\mathtt{en}$) results when the language order is reversed (6 + 6 epochs). Results are aggregated across $\mathrm{L}_{2}$ ($\mathtt{de}$ and $\mathtt{fi}$). Top: PPL per character on the $\mathrm{L}_{1}$ ($\mathtt{en}$) validation set during training. Middle: Accuracy on BLiMP during training. Bottom: Performance on GLUE at the end of training on $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$.
We illustrate the results for this experiment in Fig. 2. In general, the learning patterns through epochs are similar across conditions (except for interleaved), with variations observed mostly in the final performance. As expected, the monolingual performance is the best among all conditions. We notice that the interleaved condition performs slightly worse than the sequential condition, although the difference is more noticeable in RoBERTa models than in GPT-2 models. Both achieve lower scores than the monolingual training (the baseline for native-level language proficiency). Results from the sequential-interleaved condition differ based on the model architecture. On GPT-2 it is worse compared to sequential results (i.e., keeping $\mathrm{L}_{1}$ exposure hinders $\mathrm{L}_{2}$ learning), while on RoBERTa it is better (i.e., keeping $\mathrm{L}_{1}$ exposure helps $\mathrm{L}_{2}$ learning). Lastly, the sequential-ewc condition shows clear differences, with much worse $\mathrm{L}_{2}$ proficiency. Unsurprisingly, the reduction in plasticity through regularization has a significant (negative) impact on $\mathrm{L}_{2}$’s learning outcome.
@@
Figure 4: Summary of the $\mathrm{L}_{2}$ ($\mathtt{en}$) evaluation results for the convergence training (48 epochs). Results are aggregated across $\mathrm{L}_{1}$ ($\mathtt{de}$ and $\mathtt{fi}$). Top left: PPL per character on the $\mathrm{L}_{2}$ ($\mathtt{en}$) validation set during training on $\mathrm{L}_{2}$. Top right: PPL per character on the $\mathrm{L}_{1}$ ($\mathtt{de}$, $\mathtt{fi}$) validation set during training. Bottom left: Accuracy on BLiMP during $\mathrm{L}_{2}$ training. Bottom right: Performance on GLUE at the end of $\mathrm{L}_{2}$ training.
@@
[…]
@@
However, as we have seen from Fig. 2, $\mathrm{L}_{2}$ learning is also harmed in this case. To explore this trade-off, we vary the $\lambda$ values and test the models’ performance on both $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ at the end of training. The results are illustrated in Fig. 5. We see that the regularization strength $\lambda$ highly influences both $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ learning outcomes. When high $\lambda$ values are used (strong EWC regularization), $\mathrm{L}_{1}$ knowledge can be predictably maintained at the initial levels. However, $\mathrm{L}_{2}$ learning will be almost completely impaired. On the other side, when lower $\lambda$ values are used (weak EWC regularization), $\mathrm{L}_{2}$ learning is affected less, but $\mathrm{L}_{1}$ will not be preserved. It is also noticeable that in this case, there is a higher variance on the final $\mathrm{L}_{1}$ outcomes. As mentioned in § 5, we have selected the $\lambda$ values for all our experiments as the point where $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ performance is roughly equivalent (intersection lines are marked on the plots). Thus, we do not favor either very strong or very weak regularization, and also to match the behavior exposed by the interleaved models.
Figure 6: $\mathrm{L}_{2}$ ($\mathtt{en}$) results for each language pair at the end of training (6 epochs). Top: PPL per character on $\mathrm{L}_{2}$ ($\mathtt{en}$) at the end of training minus PPL per character from the monolingual condition. Bottom: Accuracy on BLiMP at the end of training.
@@
This experiment is motivated by the observation that CP effects can become stronger with a later age of exposure. We extend the training time for each language (thus postponing the start of $\mathrm{L}_{2}$ exposure) to 48 epochs, allowing the model weights to better converge. As this is more computationally demanding, we do only one run per condition and we also omit the sequential-interleaved condition.
The results are provided in Fig. 4, which shows that the final results are more uniform across the first three conditions, especially between interleaved and sequential. We still see a substantial loss in $\mathrm{L}_{1}$ performance in the sequential condition, indicating that longer $\mathrm{L}_{1}$ training does not lead to entrenchment, though these PPL scores are for $\mathtt{fi}$ and $\mathtt{de}$, and thus are not directly comparable to the $\mathtt{en}$ from other experiments. The performance of the interleaved setting shows a slight improvement compared to regular training, indicating that a longer training duration was beneficial (and necessary for convergence). The learning pattern for the sequential-ewc condition does not change, i.e., the final $\mathrm{L}_{2}$ performance does not significantly improve with additional training. It also appears the EWC regularization does not simply slow down $\mathrm{L}_{2}$ learning, but rather acts as a lower bound: $\mathrm{L}_{2}$ knowledge can never improve past a certain point for a given choice of $\lambda$.
@@
This experiment both addresses RQ 4 and provides more diversity in the results; the latter is especially important considering the limited selection of languages for the previous experiments. The main focus here is to find whether the choice of languages or more concretely, the degree of similarity between $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$, has any impact on the CP. For this, we consider a wider range of languages based on their relatedness to English (see § 5). To accommodate for the increase in computational demand, we consider only the GPT-2 model architecture (with configuration $\mathbf{C_{5}}$), we reduce the size of the training data to 100 million tokens, we only train with English as $\mathrm{L}_{2}$, and we only run PPL and BLiMP evaluations.
We present the findings in Fig. 6. Considering $\mathrm{L}_{2}$ PPL, we find the expected pattern, i.e., PPL generally increases (gets worse) when $\mathrm{L}_{1}$ is less closely related to it. However, the effect size is quite small (hence we plot the difference with respect to monolingual condition): only 0.1-0.3 in most cases, while the total PPL ranges from about 2.3-2.5. However, the BLiMP results do not support the same conclusion. Mostly, BLiMP scores vary seemingly at random, and indeed these differences could reflect random noise due to model initialization. As expected, $\mathrm{L}_{2}$ performance on BLiMP is greatest when $\mathrm{L}_{1}$ and $\mathrm{L}_{2}$ are just different corpora of English. But curiously, Java pretraining aids BLiMP performance more than most natural languages. The only condition where we can make the strongest case for an effect is sequential-EWC, which suggests that perhaps relatedness effects exist at the beginning of transfer but are wiped out by extensive $\mathrm{L}_{2}$ training.